Booster - Bill, MI increase intensity to 2 more turns, 1 more coming. BNB intensity start to be affected.
NuMI at 4+6 since shutdown (2+6, and sometimes 4+6), shielding assesment is limit of BNB POT. Will be loss limited after that in Booster.
Bunch rotation 1ns less . 2e17 POT is loss limit regardless of shielding.

BNB - Tom, replaced something on BNB, want to monitor system. No plans on BNB except for shutdown.

ANNIE - platform in place waiting for tank. MITPC will be limited when that comes in. Passed pORC for 60 PMTs in tank, and paddles.

MITPC - running.

MB - 

uB - Mike: software PMT trigger came online. Removing filter material (1/3 left), PMT rate reduction from 250 to 170 kHz. Will inform PMT triggers. Talk with Steve & Duane that if uB is down, MB is still up, so keep beam coming.

uB beam timing upgrade - Tia: optical devices under approval to order. Will present to uB.